OpenStudy (anonymous):

how can you use the change of base formula without a calculator for log2 74

OpenStudy (anonymous):

i have a ti-84 and dont know how to use it

OpenStudy (anonymous):

\[\log_{b}a= (\log a/\log b)\] example: \[\log_{4}6= \log 6/\log 4\]

OpenStudy (anonymous):

these dont make no sense to me

OpenStudy (anonymous):

Use the example...except use your numbers...